### v3

# Dissipation-Driven Transition of Particles from Dispersive to Flat Bands

Abstract
Flat bands (FBs) play a crucial role in condensed matter physics, offering an ideal platform to study strong correlation effects and enabling applications in diffraction-free photonics and quantum devices. However, the study and application of FB properties are susceptible to interference from dispersive bands. Here, we explore the impact of bond dissipation on systems hosting both flat and dispersive bands by calculating the steady-state density matrix. We demonstrate that bond dissipation can drive particles from dispersive bands into FBs and establish the general conditions for this phenomenon to occur. Our results demonstrate that dissipation can facilitate FB preparation, property measurement, and utilization. This opens a new avenue for exploring FB physics in open quantum systems, with potential implications for strongly correlated physics.
Introduction.— Flat band (FB) systems have attracted widespread attention due to their quenched kinetic energy, leading to eigenmodes that are compactly localized in space. This makes the system highly sensitive to interactions, often giving rise to strongly correlated phenomena such as fractional Chern insulators Chern1 ; Chern2 ; Chern3 , superconductors Chern3 ; SC1 ; SC2 ; SC3 ; SC4 , Mott insulators MI1 ; MI2 , and others. Even in non-interacting systems, FB systems can exhibit many interesting phenomena, such as the inverse Anderson transition AT1 ; AT2 ; AT3 , multifractal behavior and unconventional mobility edges ME1 ; ME2 ; ME3 , influenced by disorder or quasi-periodic potentials. FBs have also garnered significant experimental interest and have been realized in various systems, including solid-state systems exp1 ; exp2 , cavity polaritons exp3 ; exp4 ; stub2 , photonic waveguides exp5 ; exp6 ; exp7 , superconducting wire networks SCwire1 ; SCwire2 and ultracold atoms in optical lattices cold1 ; cold2 ; cold3 ; cold4 ; cold5 ; cold6 ; cold7 ; cold8 . In particular, recent advances in Moiré materials provide an intrinsic platform for studying the connection between FBs and quantum geometry, as well as the applications of FB systems Chern3 ; SC1 ; SC2 ; Moire1 ; Moire2 ; Moire3 ; Moire4 ; Moire5 .
The formation of FBs typically arises from the destructive interference of wavefunctions in specific lattice structures BAV ; BAV2 ; BAV3 , such as Lieb Lieb1 ; Lieb2 , diamond diamond1 ; diamond2 , stub stub2 ; stub1 , Kagome Kagome1 ; Kagome2 , or sawtooth Sawtooth1 ; Sawtooth2 lattices. This destructive interference leads to the formation of compact localized states (CLSs) in FBs. Given a single CLS, the whole CLS set can be generated through lattice translations. The CLS can be classified based on the minimum number $U$ of unit cells they span CLS1 ; CLS2 . As shown in Fig. 1, the cross-stitch lattice and sawtooth lattice are examples corresponding to $U=1$ and $U=2$, respectively. Notably, it is rare to find a system where all the bands are flat. Typically, dispersive (non-flat) bands are still present, which may significantly influence both experimental measurements and practical applications of FB properties. Consequently, eliminating the influence of particles in the dispersive bands has both theoretical and practical significance.
Dissipation occurs widely in various systems and profoundly impacts the properties of these systems. Although dissipation is generally considered harmful to quantum correlations, recent progress in experimental techniques for controlling various types of dissipation technology0 ; technology1 ; technology2 ; technology3 ; technology4 ; technology5 ; technology6 ; technology7 ; technology8 ; technology9 ; technology10 ; technology11 ; technology12 ; technology13 ; technology14 ; technology15 ; technology16 ; technology18 ; technology19 ; technology20 ; technology21 ; technology21s ; technology22 ; technology23 has led to growing interest in using dissipation to control quantum states or phase transitions technology3 ; technology4 ; technology5 ; technology6 ; technology7 ; technology8 ; technology9 ; technology10 ; technology11 ; technology12 ; technology13 ; technology14 ; technology15 ; technology16 ; technology20 ; technology21 ; technology21s ; technology22 ; technology23 ; phenoma1 ; phenoma2 ; phenoma3 ; phenoma4 ; phenoma5 ; phenoma6 ; PT0 ; PT1 ; PT2 ; PT3 ; PT4 ; Loc1 ; Loc2 ; PT5 ; PT6 ; PT7 ; PT8 ; PT9 ; PT10 ; PT12 . In this Letter, we investigate a type of experimentally realizable bond dissipation applied to systems with FBs and show that it can drive the system into the FB, regardless of the initial state. This means that such dissipation can drive particles from the dispersive bands into the FB, thereby serving as a control mechanism to eliminate the influence of dispersive bands. This control effect is applicable to a variety of FB systems.
General discussion on the conditions for bond dissipation driving particles into the flatband.— The dynamical evolution of an open quantum system’s density matrix under Markov approximation is governed by the Lindblad master equation GLindblad ; HPBreuer ,
$\frac{d\rho}{dt}=\mathcal{L}[\rho]=-i[H,\rho]+\Gamma\sum_{j}(O_{j}\rho O_{j}^{\dagger}-\frac{1}{2}\{O_{j}^{\dagger}O_{j},\rho\}),$
where $\mathcal{L}$ represents the Lindbladian, with the steady-state density matrix $\rho_{ss}$ corresponding to its eigenstate with a zero eigenvalue, i.e., $\mathcal{L}[\rho_{ss}]=0$. $O_{j}$ represents the jump operator that acts on a pair of sites $j$ and $j+q$, with a site-independent dissipation strength $\Gamma$, as defined by phenoma1 ; PT10 ; PZollerB1 ; PZollerB2 ; Marcos2012 ; Yusipov ; YuchengW :
$O_{j}=(c_{j}^{{\dagger}}+ac_{j+q}^{{\dagger}})(c_{j}-ac_{j+q}),$ (1)
where $a=\pm 1$ and $q\geq 1$. This dissipation, termed bond dissipation, preserves the system’s particle number but modifying the relative phase between pairs of sites separated by a distance $q$. It can drive particles into an in-phase (out-of-phase) state, corresponding to $a=1$ ($a=-1$). Such dissipation has been proposed to be realizable in cold atom systems phenoma1 ; PT10 ; PZollerB1 ; PZollerB2 and arrays of superconducting microwave resonators Marcos2012 .
We now discuss the conditions for constructing dark states of bond dissipation described by Eq.(2) in FBs. For an eigenstate $|\Psi_{n}\rangle$, if it satisfies the condition that a set of dissipative operators $O_{j}$ annihilate it, i.e.,
$\forall j,\quad O_{j}|\Psi_{n}\rangle=0,$ (2)
then $|\Psi_{n}\rangle$ is a dark state PZollerB1 . If such a dark state $|\Psi_{n}\rangle$ exists, the steady state of the system can be a pure state, given by $\rho_{ss}=|\Psi_{n}\rangle\langle\Psi_{n}|$. In a $U$-class FB system with $L$ unit cells, there are $N=L/U$ CLSs. Any eigenstate on the FB can be constructed through the superposition of these $N$ CLSs, i.e.,
$|\Psi_{n}\rangle=\sum_{j=1}^{N}A_{j}|\phi_{\text{CLS}}^{j}\rangle,$ (3)
where $A_{j}$ denotes the complex amplitude for the $j$-th CLS $|\phi_{\text{CLS}}^{j}\rangle$. We first treat a CLS as a whole, namely, $c^{\dagger}_{j}$ ($c_{j}$) in Eq. (1) represents the creation (annihilation) operator of the $j$-th CLS. From Eqs. (1, 2, 3), it is straightforward to see that when $A_{j}$, $|\phi_{\text{CLS}}^{j}\rangle$, and $q$ satisfy the following three conditions:
$A_{j}=A_{j+\kappa},\quad|\phi_{\text{CLS}}^{j+\kappa}\rangle=a|\phi_{\text{CLS}}^{j}\rangle,\quad\text{and}\quad q=\kappa,$ (4)
where $\kappa\geq 1$ is an integer, the FB state in Eq.(3) will become a dark state that satisfies Eq.(2). In FB systems, lattice translations of a single CLS generate multiple CLSs that are spatially distinct and phase-uncorrelated, allowing independent manipulation through bond dissipation to satisfy the dark state’s phase relationship requirements. Moreover, any linear combination of these CLSs constitutes an eigenstate in the FB, allowing the coefficients $A_{j}$ to be freely tuned to satisfy dark state formation conditions. Thus, the first two conditions in Eq.(4) can be satisfied for any sign of $a$. In Eqs. (3, 4), $j$ indexes the CLS, each spanning $U$ unit cells in a $U$-class FB system. For a multi-chain system, each unit cell contains sites per chain. The third condition above, where $q$ spans $\kappa$ CLSs (i.e., $q=\kappa$), implies that on each chain this corresponds to $\kappa U$ unit cells, while requiring the bond dissipation to maintain identical form across different chains. That is, the conditions for CLSs on a FB to form a dark state under bond dissipation are:
$q=\kappa U,\quad\text{and}\quad O^{1}_{j}=O^{\alpha}_{j}(\alpha>1),$ (5)
where $O^{\alpha}_{j}$ denotes bond dissipation on the $\alpha$-th chain, which has the same form as the bond dissipation on the first chain $O^{1}_{j}$. This requirement arises because the amplitude distribution and phase relationships among lattice sites within a single CLS are fixed. The bond dissipation on one chain determines the phase relationships between its sites, which in turn fixes those on the other chains to satisfy the CLS pattern. As a result, the form of bond dissipation applied to the other chains is also constrained. We need to clarify two points: 1. Eq.(5) does not require the dissipation strength $\Gamma$ to be the same across all chains. Even applying bond dissipation on only one chain (with $\Gamma=0$ on the others) can still drive the steady state into the FB. In other words, the phase relationship between lattice sites is determined through bond dissipation on a single chain, and the occupation on other chains is subsequently fixed accordingly, while also satisfying the occupation ratio required by the CLS. 2. The requirement on $q$ in Eq. (5) can be relaxed for some specific CLS distributions (see Supplemental Material SM ).
Figure 1: (a) and (c): Circles denote lattice sites, and filled circles indicate the distribution of a CLS. (a) Cross-stitch lattice with a $U=1$ CLS configuration, with the corresponding energy spectrum shown in (b). We set the inter-cell hopping to $-1$ and the intra-cell hopping to 00. (c) Sawtooth lattice with a $U=2$ CLS configuration, with the corresponding energy spectrum shown in (d). The intra-cell hopping is set to $-\sqrt{2}$, while the inter-cell hoppings are $-\sqrt{2}$ between the upper and lower chains and $-1$ between the lower chain sites. Throughout this work, we set the onsite energies to 00 and use open boundary conditions.
When the bond dissipation follows the form prescribed in Eq. (5), the system’s steady state will always populate the FB, regardless of the initial state. Below, we will illustrate this point using the cross-stitch model and the sawtooth model as examples [Fig. 1]. Both models consist of upper and lower chains, with each unit cell containing two lattice sites. The wave function is written as $|\Psi\rangle=(\dots,\psi_{j-1},\psi_{j},\dots)$, where each $\psi_{j}$ is a two-component vector representing the amplitudes on the upper and lower sites of the $j$-th unit cell. Using the eigenvalue equation $H\Psi=E\Psi$, we obtain:
$H_{-1}\psi_{j-1}+H_{0}\psi_{j}+H_{1}\psi_{j+1}=E\psi_{j},$ (6)
where the $2\times 2$ matrices $H_{1}=H_{-1}^{\dagger}$ describe the hopping between the lattice sites of neighboring unit cells, and $H_{0}$ describes the on-site potentials and the hopping between lattice sites within the same unit cell.
Cross-stitch model.— We begin by examining the impact of bond dissipation on the cross-stitch model [Fig. 1(a)]. In the absence of on-site potentials, the system features exactly one FB and one dispersive band CLS1 . We set the inter-cell hopping strength to $t_{1}$, which applies both along the same chain and between the upper and lower chains across unit cells, and set the intra-cell hopping strength to $t_{0}$ for hopping between the upper and lower chains within the same unit cell. Consequently, $H_{0}$ and $H_{1}$ in Eq.(6) are expressed as
$H_{0}=-t_{0}\begin{pmatrix}0&1\\ 1&0\end{pmatrix},\quad\quad H_{1}=-t_{1}\begin{pmatrix}1&1\\ 1&1\end{pmatrix}.$ (7)
The position of the FB is determined by the intra-cell hopping strength $t_{0}$ and is given by $E_{FB}=t_{0}$. Without loss of generality, we set $t_{1}=1$ and $t_{0}=0$ [Fig. 1(a)], such that a FB exists at $E_{FB}=0$ [Fig. 1(b)]. This FB supports a series of CLSs with the wave function $|\phi_{\text{CLS}}\rangle=(1,-1)^{T}$ [Fig. 1(a)], which is strictly localized within a single unit cell.
Figure 2: Steady-state properties of the cross-stitch model after introducing bond dissipation with $q=1$ and $O_{j}^{u}=O_{j}^{l}$. The absolute values of the steady state’s density matrix elements for (a) $a=1$ and (c) $a=-1$ in the eigenbasis of $H$. The dashed lines separate the FB region in the middle from the dispersive band regions on the sides [see Fig. 1(b)]. Eigenvalue distribution of the steady-state density matrix for (b) $a=1$ and (d) $a=-1$, both having only one non-zero eigenvalue. The eigenstate corresponding to the non-zero eigenvalue is shown in real space in the respective inset. Here, we consider $20$ unit cells. We set the dissipation strength $\Gamma=1$ on each chain. The results remain unchanged even if the strengths on the two chains are unequal, or if $\Gamma=0$ on one of them.
As specified in Eq. (5), the implemented jump operators satisfy two requirements: (1) For the cross-stitch lattice, which belongs to the $U=1$ class model, the condition reduces to $q\geq 1$; (2) The bond dissipation operators are identical for the upper and lower chains, i.e., $O_{j}^{u}=O_{j}^{l}$, where $O_{j}^{u}$ (or $O_{j}^{l}$) represents the jump operator acting on the upper (or lower) chain. In the Supplemental Material SM , we discuss the physical effects arising from $O_{j}^{u}\neq O_{j}^{l}$. Our results are independent of the dissipation strength $\Gamma$, which can vary across chains, even with $\Gamma=0$ on one chain. Without loss of generality, we set the dissipation strength on each chain to $1$. We first fix $q=1$ and $a=1$. We proceed to examine the properties of the stationary solution $\rho_{ss}$ within the eigenbasis of the Hamiltonian $H$, expressed as $\rho_{mn}=\langle\Psi_{m}|\rho_{ss}|\Psi_{n}\rangle$, where $|\Psi_{m}\rangle$ and $|\Psi_{n}\rangle$ represent the eigenstates of $H$. Fig. 2(a) illustrates that the system’s steady state occupies the FB region, which is independent of the initial state. Therefore, this bond dissipation can drive particles from the dispersive band into the FB. Diagonalizing the steady-state density matrix $\rho_{ss}$ reveals its eigenvalues, as shown in Fig. 2(b), where only one non-zero eigenvalue is present, indicating that $\rho_{ss}$ corresponds to a pure state. The inset in Fig. 2(b) displays the real-space wave function distribution of this pure state, where $i=2j-1$ (or $2j$) corresponds to the upper (or lower) chain of the $j$-th unit cell, showing the out-of-phase on the upper and lower chains within the same unit cell (consistent with the distribution of the CLS in the FB, $|\phi_{\text{CLS}}\rangle=(1,-1)^{T}$) while maintaining the in-phase across different unit cells on the same chain, a behavior resulting from setting $a=1$ in the bond dissipation. Therefore, this pure state is constructed from an equal-weight superposition of CLSs of the cross-stitch model: $|\Psi_{\text{pure}}\rangle=\sum_{j}^{L}A|\phi^{j}_{\text{CLS}}\rangle$. When $a=-1$, the steady state of the system remains in the FB region [Fig. 2(c)], and $\rho_{ss}$ still corresponds to a pure state [Fig. 2(d)]. However, this pure state exhibits alternating phases between neighboring unit cells on the same chain, as shown in the inset of Fig. 2(d), which arises because the bond dissipation with $a=-1$ selects the out-of-phase. In this case, the pure state can be expressed in terms of CLSs as: $|\Psi_{\text{pure}}\rangle=\sum_{j}^{L}(-1)^{j}A|\phi^{j}_{\text{CLS}}\rangle$.
Figure 3: The real parts of the eigenvalues of the Liouvillian superoperator for (a) $q=1$ and (b) $q=2$. Here, we fix $a=1$ and $L=20$. (c) To explain the emergence of the fourfold steady state for $q=2$, we redefine each pair of lattice sites on a single chain of the cross-stitch model as a new unit.
When $q>1$, it is numerically straightforward to show that the steady state of the system still occupies the FB region, but the number of steady states will increase. Fig. 3(a) and Fig. 3(b) respectively display the real parts of the eigenvalues, $\text{Re}(E_{\mathcal{L}})$, of the Liouvillian superoperator $\mathcal{L}$ for $q=1$ and $q=2$ with fixed $a=1$. Here $\mathcal{L}$ is represented as a $(2L)^{2}\times(2L)^{2}$ matrix, where $L$ denotes the number of unit cells. One can see that for $q=1$ and $q=2$, there are numerically one and four values, respectively, for which $\text{Re}(E_{\mathcal{L}})\rightarrow 0$, indicating that the system has one and four steady states. To understand why four steady states appear when $q=2$, we regroup the $q$ sites on each chain into a single unit, labeled as $d$, as shown in Fig. 3(c). The bond dissipation in Eq.(1) with $q=2$ affects next-nearest-neighbor sites, meaning that sites $1$ and $2$ in different units are independently influenced. This leads to the formation of two dark states that satisfy Eq.(2): $|\Psi^{(1)}_{\text{dark}}\rangle=\sum_{d}A|\phi_{\text{CLS}}^{2d-1}\rangle$ and $|\Psi^{(2)}_{\text{dark}}\rangle=\sum_{d}B|\phi_{\text{CLS}}^{2d}\rangle$, both of which lie on the FB. These two dark states span a 2D steady-state subspace, with the corresponding density matrix having a dimension of $4$. The steady-state density matrix is given by: $\rho_{ss}=\sum_{i,j=1}^{2}C_{ij}|\Psi^{(i)}_{\text{dark}}\rangle\langle\Psi^{(j)}_{\text{dark}}|$, where the coefficients $C_{ij}$ form a positive semi-definite, Hermitian and unit trace matrix. The method of utilizing such bond dissipation with $q>1$ to prepare multiple steady states can also be easily extended to systems without FBs, potentially leading to important applications.
Sawtooth model.— We next take the sawtooth model as an example to study the effect of bond dissipation on $U=2$-type FB systems. When the onsite energies are the same (set to 00 here), and the ratio of the diagonal hopping between the upper and lower chains to the baseline hopping within the lower chain is $\sqrt{2}$ [see Fig. 1(c)], i.e., the system’s Hamiltonian is given by CLS1 ; CLS2
$H_{0}=-\begin{pmatrix}0&\sqrt{2}\\ \sqrt{2}&0\end{pmatrix},\quad\quad H_{1}=-\begin{pmatrix}0&\sqrt{2}\\ 0&1\end{pmatrix},$ (8)
this configuration results in a FB, as shown in Fig. 1(d). The FB contains a series of CLSs, each of which occupies two unit cells, given by $|\phi_{\text{CLS}}\rangle=((1,0)^{T},(1,-\sqrt{2})^{T})$ [Fig. 1(c)].
Figure 4: The steady-state properties of the sawtooth model under bond dissipation with $q=2$ and $O_{j}^{u}=O_{j}^{l}$. The steady-state density matrix elements for (a) $a=1$ and (b) $a=-1$ in the eigenbasis of $H$. The dashed lines mark the boundary between the FB and dispersive band regions [see Fig. 1(d)]. Eigenvalue distribution of the steady-state density matrix for (c) $a=1$ and (d) $a=-1$, both featuring a single nonzero eigenvalue. The real-space distribution of the corresponding eigenstate is shown in the insets. Here, we fix $L=20$.
Since the sawtooth model belongs to the $U=2$ class, according to Eq. (5), the form of the applied bond dissipation satisfies $q=2\kappa$. Additionally, another requirement is the same as the second condition for the bond dissipation applied to the cross-stitch model: $O_{j}^{u}=O_{j}^{l}$. We first fix $q=2$ (i.e., $\kappa=1$). Figures 4(a) and 4(b) display the steady-state density matrix in the Hamiltonian’s eigenbasis for $a=1$ and $a=-1$, respectively, demonstrating that the system’s steady state always occupies the FB region. Moreover, in both cases, the steady-state density matrix $\rho_{ss}$ exhibits only one nonzero eigenvalue [see Figs. 4(c) and 4(d)], indicating that the final steady state is a pure state. For $a=1$, the pure state exhibits an in-phase distribution between next-nearest-neighbor lattice sites on the same chain (see inset in Fig. 4(c)), while for $a=-1$, it shows out-of-phase distribution (see inset in Fig. 4(d)). Combining with the form of the CLS in this model, these two pure states can be respectively expressed in terms of CLS as:$|\Psi_{\text{pure}}\rangle=\sum_{j}^{N}A|\phi^{j}_{\text{CLS}}\rangle$ and $|\Psi_{\text{pure}}\rangle=\sum_{j}^{N}(-1)^{j}A|\phi^{j}_{\text{CLS}}\rangle$, where $N=L/U$ represents the total number of CLS.
Conclusion and Discussion.— We have demonstrated that bond dissipation satisfying the form of Eq. (5) can drive particles from dispersive bands into FBs. Upon reaching the steady state, the dissipation can be removed, after which the density matrix evolves in time as $\rho(t)=\sum_{nm}e^{i(E_{n}-E_{m})t}\rho_{nm}|\Psi_{n}\rangle\langle\Psi_{m}|$, where $E_{n}$ and $E_{m}$ are the eigenvalues corresponding to the eigenstates $|\Psi_{n}\rangle$ and $|\Psi_{m}\rangle$ of the Hamiltonian $H$, respectively. Clearly, the diagonal elements $\rho_{nn}(t)=\rho_{nn}$ remain constant over time. This indicates that the particles occupy only the FB regions, just as in the steady state. Thus, by temporarily applying dissipation and then removing it, the Hamiltonian’s parameters and spectral structure stay unchanged, but particles originally in dispersive bands are driven into the FBs.
Our results can be experimentally investigated using cold atoms in optical lattices. Numerous models featuring FBs have already been realized in cold-atom systems cold1 ; cold2 ; cold3 ; cold4 ; cold5 ; cold6 ; cold7 ; cold8 , including the cross-stitch lattice cold3 and sawtooth lattice cold2 discussed in this work. Furthermore, the experimental scheme for implementing bond dissipation [Eq.(1)] was originally proposed using cold-atom platforms phenoma1 ; PZollerB1 ; PZollerB2 . The initial approach involved introducing a driving laser to achieve the $q=1$ case by adjusting the relationship between the wavelength of the driving laser and the optical lattice phenoma1 ; PZollerB1 ; PZollerB2 . Recent work has demonstrated that by modulating the polarization of the driving laser to achieve coupling between specific internal states, cases such as $q=2$ or even $q>2$ can be realized PT10 . While we have primarily used the cross-stitch lattice and sawtooth lattice as examples in this work, our results can be generally applied to other quasi-one-dimensional FB systems. In future research, we will extend our results to interacting systems and two-dimensional FB systems.
Using dissipation eliminates the influence of particles in dispersive bands, which facilitates the study of phenomena related to FB physics and holds potential applications in strongly correlated systems, including theoretical studies, experimental investigations, and material preparation. Our work opens a new avenue for leveraging dissipation to drive particles into FBs, offering fresh insights into the study of FBs in open systems. Although experimental realization of this studied bond dissipation in materials remains challenging, leveraging the unique properties of FBs may enable identification of feasible dissipative configurations in material systems that can generate analogous physical effects.
Acknowledgements.
This work is supported by National Key R&D Program of China under Grant No.2022YFA1405800, the Key-Area Research and Development Program of Guangdong Province (Grant No.2018B030326001), Guangdong Provincial Key Laboratory(Grant No.2019B121203002). C. Y. acknowledges support from the Key-Area Research and Development Program of Guangdong Province Grant No.2020B0303010001, Grant No.2019ZT08X324, No.2019CX01X042, No.2019B121203002.
Supplementary Material:
Dissipation-Driven Transition of Particles from Dispersive to Flat Bands
In the main text, we require identical forms of bond dissipation for both the upper and lower chains, i.e., $Q_{j}^{u}=Q_{j}^{l}$, while allowing the dissipation strengths $\Gamma$ of the two chains to differ. In the Supplementary Materials, we examine the physical consequences arising from distinct bond dissipation configurations between the upper and lower chains. We fix the dissipation strength $\Gamma$ on both chains to be $1$. We denote the elements in the bond dissipation operator $O^{u}$ applied to the upper chain as $a^{u}$ and $q^{u}$, and those in the bond dissipation operator $O^{l}$ applied to the lower chain as $a^{l}$ and $q^{l}$.
I. $a^{u}\neq a^{l}$
We take the cross-stitch model as an example, keeping $q$ of both the upper and lower chains fixed at $1$. However, the bond dissipation added to the upper chain is set to $a=1$, while the bond dissipation added to the lower chain is set to $a=-1$. Fig.S1 demonstrates that the system’s steady state does not exclusively occupy the flat band (FB) region. This can be understood as follows: While different compact localized states (CLSs) may acquire arbitrary relative phases, the intra-CLS phase difference between upper and lower chains is fixed. For instance, the CLS in the cross-stitch model takes the form $|\phi_{\text{CLS}}\rangle=(1,-1)^{T}$, indicating opposite phases on the two chains. The condition $a^{u}=1$ requires identical phases between neighboring sites on the upper chain for steady-state formation. If the steady state were composed of CLS superpositions, this would further impose identical phases between lower-chain neighbors (as all lower-chain occupations must maintain $\pi$-phase opposition to the upper chain). However, $a^{l}=-1$ favors anti-phase configurations between neighboring sites, thereby disrupting the CLS construction. Consequently, the asymmetry condition $a^{u}\neq a^{l}$ prevents complete FB occupation in the final steady state.
Figure S1: The absolute values of the steady state’s density matrix elements in the eigenbasis of $H$. We employ identical conditions and parameters as in Fig. 2 of the main text, except for modifying the previously equal chain coefficients to asymmetric values $a^{u}=1$ and $a^{l}=-1$.
II. $q^{u}\neq q^{l}$
We take the sawtooth model as an example, keeping $a$ of both the upper and lower chains fixed at $1$. However, $q^{u}$ and $q^{l}$ are different, with $q^{u}=1$ and $q^{l}=2$. Fig. S2(a)(b) and Fig. 4(a)(c) in the main text are exactly the same. This system’s steady state is occupied in the FB region and is a pure state, which also exhibits an in-phase distribution between next-nearest-neighbor lattice sites on the lower chain, while on the upper chain, neighboring lattice sites share the same phase distribution. Such a choice of bond dissipation depends on the configuration of the CLS in the sawtooth lattice. Each CLS occupies two unit cells, where the distribution within a CLS is such that neighboring lattice sites on the upper chain have the same phase (thus, when $q^{u}=1$, only $a=1$ is allowed, and $a=-1$ is not allowed), while on the lower chain, one site is occupied while the other remains unoccupied (see Fig.1(c) in the main text). Therefore, when the specific configuration of the CLS is known in advance, the form of bond dissipation can be designed based on the characteristics of the CLS, and it does not necessarily need to satisfy Eq. (5) in the main text. However, Eq. (5) in the main text is always capable of achieving a steady-state occupation in the FB, regardless of the model or the form of the CLS.
Figure S2: We adopt the same conditions and parameters as those in Fig. 4(a) of the main text, except that we modify the originally equal chain coefficients from $q^{u}=q^{l}=2$ to $q^{u}=1$ and $q^{l}=2$. Note that in this case, we fix $a=1$. $a=-1$ cannot achieve a steady-state occupation in the FB.